# DualLQR: Efficient Grasping of Oscillating Apples using Task Parameterized Learning from Demonstration

### v1

@@ V Discussion
DualLQR increased the final approach accuracy by 60% and reduced the distance travelled compared to the InfLQR. This was due to the use of the finite horizon LQR and reducing the computations in the real time loop, allowing for faster responses to changes in the target pose. In addition, the difference in required precision between the reference frames was used twice. First in the fitting of each LQR, where higher precision resulted in larger control outputs. Next, it was used again when combining the outputs. This increased the effect.
@@
In order to evaluate the performance of the methods, we used the final approach accuracy as a metric. Related work primarily used the accuracy at the final pose [21, 3]. However, this does not take the approach into account, where collisions can result in additional movement of a target or damage to a target. However, the final-approach-accuracy metric also leaves room for improvement. In the apple grasping experiment, the DualLQR was able to grasp apples successfully despite not meeting the required accuracy. The metric can be adjusted to calculate deviation from a desired path. However, a desired path would need to be defined.
@@
In the apple grasping experiment, the OptiTrack system and using a real manipulator added roughly 45 ms of latency to the control loop between a movement of the target in the world and an action of the manipulator. Combined with the dynamics required to move the manipulator, this resulted in a reduced accuracy compared to the simulation experiment. Using a camera and processing to obtain the pose of the fruit will add even more time. One of the ways this latency can be addressed is by using a predictor to estimate the future pose of the fruit, e.g. by using Model Predictive Control [22]. This can improve the performance of the proposed method in the apple grasping experiment, where the apple was oscillating consistently. However, in the case of a real apple tree, this prediction will be more challenging, due to variable flexibility in branches, wind, and other operations in the same tree. In our experiments, we found that lower control costs result in similar performance between the simulation and real environment. In a system with increased latency, even lower control costs might be required.
@@
Through the OptiTrack system, we obtained a continuous and precise pose of the apple. In a real scenario, a vision system will be less reliable. When far away, the influence of incorrect pose estimation will be minimized. As the robot approaches the fruit, the pose estimation is expected to improve, and DualLQR can adjust for theses improvements. However, when very close to the fruit, the robot can block the view of the fruit. Therefore, further sensorization of the system might be needed to obtain the pose of the apple when very close to the fruit.
@@
In the apple grasping experiment, a reduced accuracy did not yet lead to reduced grasping success. The suction cup was continuously sucking and showed a tolerance for misalignment of around 1cm. If the apple was within the tolerance at some point during the action, the apple would be grasped. This effect increased the success rate. However, contact also occurred with greater misalignment, which caused the apple to bounce off. At low control costs, repeated bouncing could occur due to the fast movement of the manipulator after the initial bounce. While the contact itself is not expected to cause damage to real apples due to the use of soft silicon, the movement caused by the interaction could cause damage to the branch or peduncle. In addition, higher accuracy would be required if an encompassing gripper was used.
@@
For the DualLQR, we used two reference frames. The methods can be extend to include any number of reference frames. This can then be extended to e.g. obstacle avoidance, if an algorithm provides the importance of the obstacle avoidance through a covariance matrix.
VI Conclusion
@@
In this work, we presented DualLQR, a novel Learning from Demonstration method aimed at approaching oscillating targets while only tracking the oscillation near the target. Through extensive simulation testing, it was found that InfLQR barely meets the required final accuracy without oscillations and drops below the required accuracy with an oscillating target. DualLQR was found to be able to meet the required final accuracy even with high oscillations. Both methods performed best at a control cost of $0.0$. Comparing these settings, DualLQR increased the accuracy 18% compared to InfLQR, with the largest improvement being an increase of 60% with high orientation oscillations. In addition, the translated and rotated distances were reduced by 3.7% and 2.8% respectively.
@@
Further testing in a simplified apple grasping experiment showed that DualLQR can successfully grasp oscillating apples, with a success rate of 99%. The optimal control cost was found to be $-1.2$. Here, grasping the apple was done in 14.9 seconds on average, compared to 15.6 seconds for the slowest control cost setting. In addition, the fruit was not touched before grasping. Higher control costs reduced the distance travelled, but also increased the time to grasp.

### v3

@@ 6 Discussion
DualLQR increased the final approach accuracy and reduced the distance travelled compared to both InfLQR and using a single LQR. Compared to the InfLQR, this was due to the use of the finite horizon LQR and reducing the computations in the real time loop, allowing for faster responses to changes in the target pose. In addition, the difference in required precision between the reference frames was used twice. First, in the fitting of each LQR, where higher precision resulted in larger control outputs. Next, it was used again when combining the outputs. By using the difference in required precision twice, more emphasis was placed on the frame with higher required precision. Compared to the single LQR, the DualLQR also increased the accuracy, while the opposite was expected. The additional reference frame of the DualLQR will influence the movement even during the final approach. Most noticeably, the DualLQR had less variation. This indicates that the additional reference frame dampened the control output, resulting in the more accurate approach.
@@
In order to evaluate the performance of the methods, we used the final approach accuracy as a metric. Related work primarily used the accuracy of the final pose [13, 19]. However, this does not take the approach into account, where collisions can result in additional movement of a target or damage to a target. However, the final-approach-accuracy metric also leaves room for improvement. In the apple grasping experiment, the DualLQR was able to grasp apples successfully despite not meeting the required accuracy. The metric can be adjusted to calculate deviation from a desired path. However, a desired path would need to be defined.
The experiments provided an empirical evaluation of the system robustness. There are systemic limitations to the system’s robustness because of workspace and acceleration limitations of the manipulator. The acceleration limits of the fruit are likely lower, as the fruit would become detached otherwise. The selection of target fruits needs to consider where in the workspace the fruit is hanging to ensure the fruit will stay within the workspace even under oscillations.
In the simulation experiment, the oscillation did not decay, while this would have been more realistic and in line with the apple grasping experiment.
@@
In the apple grasping experiment, the OptiTrack system and manipulator added roughly 45 ms of latency to the control loop between a movement of the target in the world and an action of the manipulator. Combined with the dynamics required to move the manipulator, this resulted in a reduced accuracy compared to the simulation experiment. Using a camera and processing to obtain the pose of the fruit will add even more time. This latency could be addressed by using a predictor to estimate the future pose of the fruit, e.g. by using Model Predictive Control [9]. This can improve the performance of the proposed method in the apple grasping experiment, where the apple was oscillating consistently. However, in the case of a real apple tree, this prediction will be more challenging, due to variable flexibility in branches, wind, and other operations in the same tree. Additionally, it would increase the latency even more, which reduces its applicability. In our experiments, we found that lower control costs result in similar performance between the simulation and real environment. In a system with increased latency, even lower control costs might be required.
@@
Through the OptiTrack system, we obtained a continuous and precise pose of the apple. In a real scenario, a vision system will be less reliable. When far away, the influence of incorrect pose estimation will be minimized. As the robot approaches the fruit, the pose estimation is expected to improve, and DualLQR can adjust for these improvements. However, when very close to the fruit, the robot can block the view of the fruit. Therefore, further sensorization of the system might be needed to obtain the pose of the apple when close to the fruit.
The grasp of the apples was detected based on the pressure value in the suction cup. While no false positives occurred during our experiments, additional sensors can be implemented to provide a second measurement of fruit pressure. These could use a FT-sensor to detect the weight of the fruit in the suction cup, or measure the deformation of the suction cup. Alternatively, as long as the false positive does not result in damage, the pick can be attempted again.
@@
In the apple grasping experiment, a reduced accuracy did not yet lead to reduced grasping success. The suction cup was continuously sucking and showed a tolerance for misalignment of around 1 cm. If the apple was within the tolerance at some point during the action, the apple would be grasped. This effect increased the success rate. However, contact also occurred with greater misalignment, which caused the apple to bounce off. At low control costs, repeated bouncing could occur due to the fast movement of the manipulator after the initial bounce. While the contact itself is not expected to cause damage to real apples due to the use of soft silicon gripper, the movement caused by the interaction could cause damage to the branch or peduncle. In addition, higher accuracy would be required if an encompassing gripper was used.
@@
For the DualLQR, we used two reference frames. The method can be extended to include any number of reference frames. This can even be extended to include for example obstacle avoidance strategies, if an algorithm provides the importance of the obstacle avoidance through a covariance matrix.
To facilitate the reproducibility of this work, we made the code to use the proposed method available, including the set-up for the simulation and apple grasping experiment. The datasets and trained models are available upon request. Most of the set-up consisted of off-the-shelf components. The end-effector was custom-made, for which the 3D models can be made available upon request. As the performance of the model depends on the demonstrations provided, training the models with a new set of demonstrations will influence the achievable performance.
7 Conclusion
@@
In this work, we presented DualLQR, a novel Learning from Demonstration method aimed at approaching oscillating targets while only tracking the oscillation near the target. Through extensive simulation testing, it was found that InfLQR barely meets the required final accuracy without oscillations and drops below the required accuracy with an oscillating target. On the other hand, a single LQR reached the required accuracy, but required a large increase in the distance travelled to reach the target. DualLQR was found to meet the required final accuracy even with high oscillations, while travelling the least distance. Compared to InfLQR, DualLQR increased the accuracy by 18%, reduced the translated distance by 3.7% and the rotated distance by 2.8%. Compared to a single LQR, DualLQR increased the accuracy by 7.0%, reduced the translated distance by 26% and the rotated distance by 6.8%.
@@
Further testing in a simplified apple grasping experiment showed that DualLQR can successfully grasp oscillating apples, with a success rate of 99%. The optimal control cost was found to be $-1.2$. Here, grasping the apple was done in 14.9 seconds on average, compared to 15.6 seconds for the slowest control cost setting. In addition, the fruit was not touched before grasping.
{credits}
7.0.1 Acknowledgements
This work was partially funded by the Dutch Ministry of LVVN (Agriculture, Fisheries, Food Security, and Nature), under the project code KB-38-001-005, and by the Netherlands Organisation for Scientific Research (NWO grant 17626, project Synergia).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
7.0.3 Electronic supplementary material
A supporting video can be found at https://youtu.be/2bY84MN53tA